# Trajectory First: A Curriculum for Discovering Diverse Policies

### v1

@@ 5 Experiments
Environments. We conduct our experiments focusing on environments from robotics. We use two environments for evaluation: (a) a cylinder avoidance task where a rod attached to a robot gripper must be navigated through a maze without collisions (see Fig. 2). The agent is rewarded for minimizing the distance to the goal line and penalized for touching obstacles. This task enables a rich set of behaviors since many paths through the maze are possible. To avoid the trivial solution of moving over the top of all obstacles, we fix the $z$-position of the rod and use $xy$-endeffector position control. We train 10 different skills for this task. (b) Second, we use a cube pushing task in which a 3d-pointmass robot is tasked to push a cube as far away from the center of a table as possible. This task enables a rich set of behaviors since there are no contact encouraging terms in the reward, so any object manipulation is conceivable. We train 4 skills for this environment. For full environment details, we refer to Appendix B.
Baselines. We compare our method to two baselines: First, we use plain diversity optimization with optimality constraints, without additional exploration, to investigate how well these methods explore. Second, we use Gaussian action noise $\epsilon\sim{\cal N}(0,I)$ as a non-diversity-specific exploration method. We implement the baseline diversity optimization as described in Domino [11], and base our code on the public implementation thereof [70]. For both baselines, we do not estimate successor features and use the ground truth observations as state features to guarantee maximum possible performance [11, 71]. As stated above, all code is based on SAC [21], and uses observation normalization, critic ensembling, and layer norm. For further implementation details, we again refer to Appendix B.
@@
To answer this question, we look at the performance of running diversity optimization without any additional exploration mechanism. We observe in Fig. 2 that the diversity optimization algorithm Domino underexplores the domain, which leads to little behavioral diversity. While the method successfully solves the problem, only a few behavioral modes are found. For the cube task, we see that multiple distinct behaviors can be found, but some skills overlap (Fig. 3). This highlights that learning a set of diverse skills requires the discovery of a sufficient number of behaviors in the first place. Further, our quantitative results (Fig. 4) highlight that online exploration only partly resolves this issue, as the overall diversity of the solutions is still limited, which we can also see in Fig. 2.
@@
To answer this question, we consider the results of training agents using the proposed curriculum. As we can see in Fig. 2, using the curriculum enables to discover almost all paths though the maze. Similarly, in Fig. 3, the curriculum exploration is the only method that finds pushes along four different directions of the table. All other methods tend to explore more subtle variations such as the final orientation of the cube once it is pushed off the table. These qualitative findings are corroborated quantitatively in Fig. 4, which shows that the proposed curriculum clearly produces the most diverse policies.
@@
This question aims to investigate whether the diversity objective during RL is needed given diverse data from the first stage of the curriculum. To answer this question, we run constraint-free population-based training [72, PBT] based on the CNS data, but only maximizing policy return. Our results in Fig. 5a show that using the diversity objective indeed increases the policy diversity. At the same time, we observe a slight decrease in task performance when optimizing for diversity, which we explain by the objective from Eq. 1, which permits a certain amount of slack $\alpha$. We believe that these observations provide valuable insights into RL finetuning for diversity. As stated above, prior finetuning literature neglected diversity, and we believe that the presented recipe closes this gap. Further, we investigate whether the inclusion of the Lagrange multiplier in the CNS formulation improves the evolutionary trajectory optimization in comparison to the scalarized novelty search formulation from previous work [10]. Fig. 5b displays the population entropy and population reward across novelty search iterations. We observe that while there are no differences in rewards, the population entropy is higher for the proposed CNS objective. This demonstrates that the generalized formulation of novelty search may be a valuable tool for exploration and discovery in the future.
@@ 6 Conclusion & Limitations
We introduced a two-stage, trajectory-first curriculum that can discovering diverse skills in challenging robotic domains. In the first phase, we use a constrained novelty search evolution strategy to explore trajectories. In the second stage, we train a set of diverse reactive control policies given the CNS data. We have shown in our experiments that: (a) naïvely applying constrained diversity objectives in policy space leads to under-exploration and thus fails to discover truly diverse skills (Fig. 2). (b) By first exploring diverse trajectories using constrained novelty search, the diversity optimization can be improved (Fig. 4). (c) Using proper performance constraints during evolutionary novelty search improves sample set entropy over scalar formulations (Fig. 5a). (d) Using diverse data for policy training alone does not guarantee truly diverse policies, but diversity objectives are still required to maintain full diversity (Fig. 5b).
Despite these advances, our approach is not without limitations. We observe a high variance in task returns (Fig. 4), which we hope to alleviate in the future. For instance, it may be possible to improve the data-based finetuning part of our pipeline, by simply increasing the number of seeds in our evaluation. Moreover, despite providing valuable insights, we aim to further understand how exploration can improve diversity optimization. We plan to compare to additional baselines such as data-based exploration methods [27, 55]. Finally, our method introduces new hyperparameters to the diversity optimization. In the future, we aim to conduct sensitivity analyses to gain a more profound understanding of the approach’s hyperparameters.
Acknowledgments and Disclosure of Funding
This research was funded by the Amazon Development Center Germany GmbH.
@@ B.1 Environments
where $s_{x}$ denotes the $x$ position of the rod, $target_{x}$ is the coordinate of the goal line, while $\mathds{1}_{collision}(s)$ is a collision checking function. Further $x_{max}$ is the maximum $x$ coordinate that is admissible, which we use as offset to guarantee position rewards. We additionally clip the rewards to be in $[0,2]$ to improve training stability. We choose $\beta=20$ for our experiments. To further simplify the task, we bound the $xy$-positions to $[-4.5,4.5]^{2}$, which we implement by clipping.
@@
Here $\lVert xyz_{cube}-xyz_{table}\rVert$ denotes the position difference between cube and table center of mass. We approximate $\dot{q}$ by first order finite differences as $\dot{q}(x)\approx x_{t+1}-x_{t}$, which we find to produce better training results than using the velocities from the Mujoco simulator that we use [73]. We choose $\beta=1/2$.
@@ B.2 Implementation
We implement all algorithms in JAX [74]. We implement Constrained Novelty Search based on the CMA-ES implementation from evosax [75]. The code for all RL agents is based on Domino [11], the public implementation thereof Grillotti et al. [70], and the STOIX ecosystem [76]. We follow Zahavy et al. [11] in the choices of all hyperparameters for Domino with exceptions detailed below. For all baselines, we use ground truth observations as state features, since they are low-dimensional and should thus yield the best performances [11, 71]. For CNS, we use the aforementioned random projections since they are an elementary part of the method. For a full list of hyperparameters, we refer to Table 1.
@@
As stated in Section 3.1 we follow design choices from prior work in using layer normalization and observation normalization. We use the same architectures for actor and critic networks, however we perform critic ensembling for regularization and use separate heads for extrinsic and intrinsic values. Our MLP backbone resembles that of Lee et al. [77], but we replace the layer norm in between block by a dynamic tanh [78]. All networks are optimized using Adam [79].
@@
We implement CNS based on the CMA-ES implementation in evosax. Before combining extrinsic and intrinsic fitness, we normalize both values within each subpopulation. Further, we use simple gradient descent to update the Lagrange multipliers. For higher optimization stability, we only update these parameters every iteration, but then perform 200 steps of gradient descent. To prevent gradient saturation due to the usage of sigmoids on the Lagrange multipliers, we bound them to make sure that they remain in a reasonable range. Similar to Domino, we also fix the first Lagrange multiplier to $1$, so we can estimate $v^{*}$ based on this population. In practice, we found it more stable, however, to choose $v^{*}=\max_{\omega\in\Omega}v(w)$. Note that we follow this choice only during CNS.
@@ Appendix C Additional Results
For completeness, we report the results of every seed for the obstacle avoidance task in Fig. 8. We see that our method learns the most diverse policy set for each seed.

### v2

@@ 5 Experiments
Environments. We conduct our experiments focusing on environments from robotics. We use two environments for evaluation: (a) a cylinder avoidance task where a rod attached to a robot gripper must be navigated through a maze without collisions (see Fig. 2). The agent is rewarded for minimizing the distance to the goal line and penalized for touching obstacles. This task enables a rich set of behaviors since many paths through the maze are possible. To avoid the trivial solution of moving over the top of all obstacles, we fix the $z$-position of the rod and use $xy$-endeffector position control. We train 10 different skills for this task. (b) Second, we use a cube pushing task in which a 3d-pointmass robot is tasked to push a cube as far away from the center of a table as possible. This task enables a rich set of behaviors since there are no contact encouraging terms in the reward, so any object manipulation is conceivable. We train 4 skills for this environment. For full environment details, we refer to Appendix˜B.
Baselines. We compare our method to two baselines: First, we use plain diversity optimization with optimality constraints, without additional exploration, to investigate how well these methods explore. Second, we use Gaussian action noise $\epsilon\sim{\cal N}(0,I)$ as a non-diversity-specific exploration method. We implement the baseline diversity optimization as described in Domino (Zahavy et al., 2023), and base our code on the public implementation thereof (Grillotti et al., 2024). For both baselines, we do not estimate successor features and use the ground truth observations as state features to guarantee maximum possible performance (Zahavy et al., 2023; G Leon et al., 2024). As stated above, all code is based on SAC (Haarnoja et al., 2018), and uses observation normalization, critic ensembling, and layer norm. For further implementation details, we again refer to Appendix˜B. For evaluation we report the mean return across all policies across 5 seeds as well as the mean diversity defined by Eq. (2), but using ground truth states of the rollout instead of expected features for diversity metric computation.
@@
To answer this question, we look at the performance of running diversity optimization without any additional exploration mechanism. We observe in Fig. 2 that the diversity optimization algorithm Domino underexplores the domain, which leads to little behavioral diversity. While the method successfully solves the problem, only a few behavioral modes are found. For the cube task, we see that multiple distinct behaviors can be found, but some skills overlap (Fig.˜3). This highlights that learning a set of diverse skills requires the discovery of a sufficient number of behaviors in the first place. Further, our quantitative results (Fig.˜4) highlight that online exploration only partly resolves this issue, as the overall diversity of the solutions is still limited, which is visible in Fig.˜2.
@@
To answer this question, we consider the results of training agents using the proposed curriculum. As we can see in Fig.˜2, using the curriculum enables to discover almost all paths though the maze. Similarly, in Fig.˜3, the curriculum exploration is the only method that finds pushes along four different directions of the table. All other methods tend to explore more subtle variations such as the final orientation of the cube once it is pushed off the table. These qualitative findings are corroborated quantitatively in Fig. 4, which shows that the proposed curriculum clearly produces the most diverse policies.
@@
This question aims to investigate whether the diversity objective during RL is needed given diverse data from the first stage of the curriculum. To answer this question, we run constraint-free population-based training (Jaderberg et al., 2017, PBT) based on the CNS data, but only maximizing policy return. Our results in Fig. 5a show that using the diversity objective indeed increases the policy diversity. At the same time, we observe a slight decrease in task performance when optimizing for diversity, which we explain by the objective from Eq.˜1, which permits a certain amount of slack $\alpha$. We believe that these observations provide valuable insights into RL finetuning for diversity. As stated above, prior finetuning literature neglected diversity, and we believe that the presented recipe closes this gap. Further, we investigate whether the inclusion of the Lagrange multiplier in the CNS formulation improves the evolutionary trajectory optimization in comparison to the scalarized novelty search formulation from previous work (Conti et al., 2018). Fig. 5b displays the population entropy and population reward across novelty search iterations. We observe that while there are no differences in rewards, the population entropy is higher for the proposed CNS objective. This demonstrates that the generalized formulation of novelty search may be a valuable tool for exploration and discovery in the future.
@@ 6 Conclusion & Limitations
We introduced a two-stage, trajectory-first curriculum that can discovering diverse skills in challenging robotic domains. In the first phase, we use a constrained novelty search evolution strategy to explore trajectories. In the second stage, we train a set of diverse reactive control policies given the CNS data. We have shown in our experiments that: (a) naïvely applying constrained diversity objectives in policy space leads to under-exploration and thus fails to discover truly diverse skills (Fig.˜2). (b) By first exploring diverse trajectories using constrained novelty search, the diversity optimization can be improved (Fig.˜4). (c) Using proper performance constraints during evolutionary novelty search improves sample set entropy over scalar formulations (Fig.˜5a). (d) Using diverse data for policy training alone does not guarantee truly diverse policies, but diversity objectives are still required to maintain full diversity (Fig.˜5b).
Despite these advances, our approach is not without limitations. We observe a high variance in task returns (Fig. 4), which we hope to alleviate in the future. For instance, it may be possible to improve the data-based finetuning part of our pipeline, by simply increasing the number of seeds in our evaluation. Moreover, despite providing valuable insights, we aim to further understand how exploration can improve diversity optimization. We plan to compare to additional baselines such as data-based exploration methods (Liu & Abbeel, 2021; Burda et al., 2019). Finally, our method introduces new hyperparameters to the diversity optimization. In the future, we aim to conduct sensitivity analyses to gain a more profound understanding of the approach’s hyperparameters.
Acknowledgements
This research was funded by the Amazon Fulfillment Technologies and Robotics team.
Supplementary Materials
The following content was not necessarily subject to peer review.
@@ B.1 Environments
where $s_{x}$ denotes the $x$ position of the rod, $target_{x}$ is the coordinate of the goal line, while $\mathds{1}_{collision}(s)$ is a collision checking function. Further $x_{max}$ is the maximum $x$ coordinate that is admissible, which we use as offset to guarantee position rewards. We additionally clip the rewards to be in [0,2][0,2][ 0 , 2 ] to improve training stability. We choose $\beta=20$ for our experiments. To further simplify the task, we bound the $xy$-positions to $[-4.5,4.5]^{2}$, which we implement by clipping.
@@
Here $\lVert xyz_{cube}-xyz_{table}\rVert$ denotes the position difference between cube and table center of mass. We approximate $\dot{q}$ by first order finite differences as $\dot{q}(x)\approx x_{t+1}-x_{t}$, which we find to produce better training results than using the velocities from the Mujoco simulator that we use (Todorov et al., 2012). We choose $\beta=1/2$.
@@ B.2 Implementation
We implement all algorithms in JAX (Bradbury et al., 2018). We implement Constrained Novelty Search based on the CMA-ES implementation from evosax (Lange, 2022). The code for all RL agents is based on Domino (Zahavy et al., 2023), the public implementation thereof Grillotti et al. (2024), and the STOIX ecosystem (Toledo, 2024). We follow Zahavy et al. (2023) in the choices of all hyperparameters for Domino with exceptions detailed below. For all baselines, we use ground truth observations as state features, since they are low-dimensional and should thus yield the best performances (Zahavy et al., 2023; G Leon et al., 2024). For CNS, we use the aforementioned random projections since they are an elementary part of the method. For a full list of hyperparameters, we refer to Table˜1.
@@
As stated in Section˜3.1 we follow design choices from prior work in using layer normalization and observation normalization. We use the same architectures for actor and critic networks, however we perform critic ensembling for regularization and use separate heads for extrinsic and intrinsic values. Our MLP backbone resembles that of Lee et al. (2024), but we replace the layer norm in between block by a dynamic tanh (Zhu et al., 2025). All networks are optimized using Adam (Kingma & Ba, 2015).
@@
We implement CNS based on the CMA-ES implementation in evosax. Before combining extrinsic and intrinsic fitness, we normalize both values within each subpopulation. Further, we use simple gradient descent to update the Lagrange multipliers. For higher optimization stability, we only update these parameters every iteration, but then perform 200 steps of gradient descent. To prevent gradient saturation due to the usage of sigmoids on the Lagrange multipliers, we bound them to make sure that they remain in a reasonable range. Similar to Domino, we also fix the first Lagrange multiplier to 111, so we can estimate $v^{*}$ based on this population. In practice, we found it more stable, however, to choose $v^{*}=\max_{\omega\in\Omega}v(w)$. Note that we follow this choice only during CNS.
@@ Appendix C Additional Results
For completeness, we report the results of every seed for the obstacle avoidance task in Fig.˜8. We see that our method learns the most diverse policy set for each seed.